OpenStudy (anonymous):

9x^3-z If x=-2 and z=8

OpenStudy (anonymous):

I got -80 is this right?

OpenStudy (anonymous):

Yup, 9(-2)^3 -8 = 9*-8 -8 = -72 -8 = -80